OpenStudy (anonymous):

evaluate integral: S (8x^4 + 5)^3 (32x^3) dx

OpenStudy (anonymous):

try u = 8x^4 + 5

OpenStudy (lalaly):

u= 8x^4 +5 du= 32x^3 dx substitute S u^3 du = u^4/4 so it is equal to (8x^4 +5)^4 /4

OpenStudy (anonymous):

got it